Compute the point advertising elasticity of demand (%ΔQX/%ΔA) of a medical supplier if the business spent P10,000 from the previous budget of P8,000 and the sales increased from P100,000 to P200,000. Interpret your answer.


1
Expert's answer
2021-03-04T10:36:41-0500

The point advertising elasticity of demand can be calculated as follows:


"E_A=\dfrac{\%\Delta Q_X}{\%\Delta A}=\dfrac{\dfrac{P200000-P100000}{P100000}}{\dfrac{P10000-P8000}{P8000}}=4.0."

As we can see from calculations, 25% increase in advertising expenditures increases demand by 4%.